OpenStudy (anonymous):

Find the measure of the three angles of a triangle, if the second is 18 degrees more than five times the first and the third is three times the measure of the first.

OpenStudy (anonymous):

A, B, C. B=5A+18 C=3A Triangle sum theorem says that all three angle measures add to 180 degrees. so A + B + C = 180 substituting, we get A + 5A + 18 + 3a = 180 and the resulting equation can be solved with simple algebra.